OpenStudy (anonymous):

what is the energy of a photon that emits a light of frequency 4.47 x 10^(14) Hz?

OpenStudy (anonymous):

Energy = hv, where v is the frequency

OpenStudy (anonymous):

how do i convert Planck's constant to Hz?

OpenStudy (jfraser):

you don't. plank's constant is a constant. it has units of J*s, so when you mulitply by frequency (\(\frac{1}{s}\)) then the units of seconds cancel

OpenStudy (anonymous):

E=hf h= 6.63 x10^-34 f = 4.47 x 10^14 E = 6.63 x 10^-34 X 4.47 x 10^14. =. 2.96 x 10^-19J